# A Framework for Guided Motion Planning

### v1

@@ IV Related work: categories of guidance
We separate guiding space methods into three main categories: robot modification, environment modification and experience based guidance. We find these subcategories natural and comprehensive, as the entire motion planning problem is encapsulated in the structure of the configuration space, which can be learned from experience, or is induced by its robot and environment constraints.
IV-A Robot Modification
@@
There are many natural ways one might modify a robot description to obtain a closely related configuration space. One of the most natural projections is to define $\mathcal{S}$ as workspace itself, ignoring the robot entirely and treating it as a fully controllable point robot in the environment [19, 13]. Some prior work ignores kinodynamic constraints on the robot [26], thereby yielding a guiding space where the robot is locally controllable. For chain-link robots, such as manipulators, a common guiding space involves focusing on a subset of the degrees of freedom [24]. Such methods often use heuristics in deciding which subsets to consider, e.g. focus on the first joints of a manipulator. Finally, some work attempts to learn a subspace of the robot’s degrees of freedom, e.g. using statistical projections such as PCA [25].
@@
All the above robot modification methods produce a search space as opposed to a guiding space. For example, workspace guidance methods often build a sparse graph in workspace, such as the medial axis [13]. Therefore there is often a secondary guided search that occurs in the modified robot space, e.g., the medial axis graph. This both highlights the hierarchical nature of guidance and helps isolate contributions - the medial axis is a useful structure for guiding search in workspace, and workspace is useful for guiding search in c-space.
IV-B Environment Modification
@@
Narrow passages are a known bottleneck for randomized sampling algorithms. As such, some works modify obstacles in an attempt to widen passages [6, 29]. Lazy planning methods [10, 18] also perform a type of environment modification by ignoring constraints during initial planning, and incorporating them later as needed. When removing constraints, paths in $\mathcal{S}$ are not necessarily valid in $\mathcal{C}$, and thus often such methods iteratively or hierarchically compute a guiding space (e.g., the initial lazy graph), use $h$ to produce samples near the unvalidated path, and then repeat with a new guiding space informed by the recently acquired samples (e.g., remove edges found to be invalid).
In contrast to the removal of environment constraints, adding constraints is less common, but can be done to simplify geometries and therefore collision detection [17]. Moreover, paths found when adding constraints are always valid in the original search space, and are thus easy to reuse. In other words, if we take $h$ to be the cost of paths in the modified space, constraint removal provides optimistic lower bounds whereas constraint addition produces conservative upper bounds.
IV-C Experience Based Guidance
The most common version of experience based guidance assumes the dataset contains a variety of environments all for the same robot. Consequently, database queries often consist of evaluating the degree to which paths in the dataset violate discovered constraints in the current motion planning instance [7]. While some methods only query using the task endpoints [12], more recent approaches propose to query (or filter) the database using local task features derived from workspace occupancy [11]. Another approach projects trajectories onto a lower dimensional abstract space via auto-encoding [20] or metrics like dynamic time warping [1], which improves database query.
Some methods perform multiple queries, where the guiding space effectively becomes the composite space of multiple independent motion planning problems from the dataset, e.g., through gaussian mixture models [11].
@@ IV-C Experience Based Guidance
In all learning based approaches a major challenge is gathering appropriate data, leading to methods which simplify the problem by learning local guidance through learned local planners [15]. Some work [5] proposes to gather appropriate data for training by using probabilistic roadmaps.
Figure 3: Example guiding spaces of each algorithm. We display the starts and goals of the rectangular robot for RRT, LazyPRM, and PathDatabase to indicate that these guiding spaces are in c-space, whereas MedialAxis is in workspace. For PathDatabase we display the paths from the database that pass near the goal.
V Evaluation metric
V-A Sampling Efficiency
Often we evaluate sampling based motion planning methods using algorithm level holistic metrics such as running time or number of samples. While minimizing such metrics is a reasonable end goal, they can be sensitive to implementation details, and focusing on them can obscure which aspect of a proposed contribution is responsible for the results. In the context of guided motion planning, we now present a new method to evaluate quality of guidance.
In this work we focus on node selection as opposed to node expansion, and note that evaluating node expansion can be reduced to evaluating a node selection procedure between the set of possible expansions at a given node.
@@ V-A Sampling Efficiency
Intuitively, we aim to measure the difference between these two distributions, and while the KL-Divergence provides such a measure, it can be dominated by small regions in space where the two distributions have largest difference (e.g., when one distribution assigns zero probability to a region in which the other is non-zero). In sampling based motion planning we often aren’t worried about a few bad samples, so long as good ones are made with high enough probability so the bounded and symmetric Jensen-Shannon Divergence metric provides a smoother comparison, i.e.,
@@
In shortest path problems we are interested in (efficiently) computing a boundedly optimal $(s,t)$ path in $\mathcal{C}$. Let $\delta_{v}$ be the suboptimality of $v\in T$, $\delta_{v}=d_{T}(s,v)+d_{\mathcal{C}}(v,t)-d_{\mathcal{C}}(s,t)>0$, where $d_{T}$ represents the distance between nodes on the tree and $d_{\mathcal{C}}$ is the true shortest path between configurations in $\mathcal{C}$. Now let $\tau_{v}$ be the work remaining for a search starting at $v$, $\tau_{v}=d_{\mathcal{C}}(v,t)$. We propose to define $Q$ with respect to these two quantities:
$Q_{T}(v)\propto\exp(-\frac{\delta_{v}}{\delta})\exp(-\frac{\tau_{v}}{\tau})$
$=\frac{e^{-(\delta_{v}/\delta+\tau_{v}/\tau)}}{Z},Z=\sum_{x}e^{-(\delta_{x}/\delta+\tau_{x}/\tau)}$
@@
V-B Axes for Evaluating Guidance
The metric described above measures the quality of each individual node expansion step of the guided search algorithm, but there are other options for evaluating guidance.
V-B1 Environment level
V-B2 Sample level
In other words, single-step evaluation considers distributions over trees $T$, whereas in multiple-sample evaluation considers distributions over $\mathcal{C}$, implicitly described by the distribution of valid trees produced by an algorithm. We refer the reader to [4] for a more in depth discussion of multiple-sample evaluation, and here summarize that in that setting our metric reduces to measuring the entropy of the observed sampling distribution. In this case lower entropy means better guidance.
Figure 5: RRT, LazyPRM, MedialAxis, and PathDatabase (Lightning) evaluations. All plots are averages over 128 random seeds and evaluation is done with a smoothed target distribution that is at least $\epsilon=0.0001$ at each node on the tree. When a seed finishes the search we consider its sampling efficiency to be 00 for all future samples (i.e., sampling the goal). Runtime in PathDatabase only includes search time and not database creation, and the guidance quality is averaged over the entire run instead of shown per iteration as in the other algorithms.
VI Refactoring Existing Algorithms
@@
Existing algorithms such as those cited in Section IV are not implemented in the form of Algorithm 1, rather each uses their respective guiding space in a different way. In the next section we conduct experiments on four guided search algorithms, RRT [23], LazyPRM [10], DR-RRT [13], and Lightning [7]. Figure 4 shows examples of trees produced by these algorithms for solving the tasks depicted in Figure 2, and Figure 3 visualizes their guiding spaces.
@@
LazyPRM builds a probabilistic roadmap without performing collision detection with the environment (or only doing so for vertices but not edges). Our implementation grows a tree using the length of the shortest path on this lazy graph as guidance for selecting which node to expand on the tree. When an invalid sample is found in c-space, we pass this information back to the guiding space (the lazy graph) by deleting nearby vertices/edges.
@@
DR-RRT uses the medial axis skeleton computed in workspace to guide the robot. The original proposed implementation uses the edges of this skeleton to define “dynamic regions” along which to sample, and does so by creating balls around these regions and projecting back into c-space. For robots whose workspace configuration is a subset of the dimensions of their c-space this is easy, but for more complex c-space the general form of such a procedure involves an inverse-kinematics projection. Our new implementation, which we call MedialAxisGuidance is more straightforward - we project c-space nodes $v$ with forward-kinematics into workspace, then project these workspace poses onto the nearest medial axis edge $e_{v}$. The corresponding guidance value is then the distance along the skeleton edges from the projected edge to the projection of the goal, $d(e_{v},e_{t})$. Whenever a node $v$ fails to expand due to obstacle collision we increase the weight of $e_{v}$, thereby reducing the chance that the guidance encourages exploration along that edge.
@@
Lightning uses a database of paths from related motion planning problems to guide exploration. Namely, given a motion planning problem, a path that was produced by the most similar task (start and goal) is selected from the database, and then copied into the current c-space. Where the path is invalid the algorithm fixes the solution by running a bi-directional RRT from the two ends of the breakage. We highlight that this sub-problem can be just as difficult as the original, and moreover that the full power of the database is never used. Our new implementation, which we call PathDatabaseGuidance uses the length of the queried path as the guidance value. As with LazyPRM, failed sample information is passed down to the guiding space by filtering out paths that pass close to invalid regions. If no node can locally plan to any path, then we default to RRT until a new node is found from which there is a visible path. We note that this implementation is relatively similar to [1], but we query the database using task similarity as in [7] rather than dynamic time warping, and more importantly we actively filter the database using information from the search tree.
VII Experiments
@@
We design environments intended to demonstrate both pros and cons of each guiding space mentioned in Section VI, using a rectangular mobile base (3-dimensional c-space) in a 2-dimensional workspace.
@@
Next we test PathDatabaseGuidance on a dataset of RandomSimplePassages, an environment with a randomly placed narrow passage between two rooms. We show how guidance improves as dataset size increases, while runtime gets worse (as querying the database takes longer), showing how such traditional metrics can obscure properties of algorithms because they are implementation dependent. This matches the intuition mentioned in Section III-C regarding the trade-off between guidance quality and the difficulty of computing it.
VII-A Multi-guidance search
Finally, we create a hybrid environment TrapCup which combines features form both the Trap and Cup environments. We now design a hybrid guiding space that is a combination of LazyPRM and MedialAxis. In other words, it creates both guiding spaces (i.e., the cross product) and uses a rule to determine how to select or combine the heuristics from the two. In the experiments shown in Figure 6 we simply swap the source of the guidance each time the search collides with an obstacle. Intuitively, each form of guidance faces certain local minima and by using multiple guiding spaces we can quickly escape those minima by passing control to another guiding space. Notice that the individual guidance plots for each guiding space indicate that these local minima are distinct (e.g., the curves have different peaks), and thus we see the hybrid guiding space out-performs its parts. Moreover, using number of samples alone as a metric would obscure this fact since MedialAxis and LazyPRM take about the same number of samples to find the goal. These results parallel what one would expect from running a multi-heuristic A* search [2] as opposed to single heuristic search, and highlights the value of framing sampling based motion planning as guided search; we now have a structured way of combining different algorithmic ideas from the literature as multiple guiding spaces, whereas the original implementation are not so conducive to modularization.
VIII Conclusion

### v2

@@ 4 Categories of Guidance
We separate guiding space methods into three main categories: robot modification, environment modification and experience-based guidance. We find these subcategories natural and comprehensive, as the entire motion planning problem is encapsulated in the structure of the configuration space, which can be learned from experience, or is induced by its robot and environment constraints.
In this section, as we define the categories of guidance, we discuss how existing work fits into each category. Note that much of the work cited below does not explicitly describe guidance as a contribution, yet here we identify that they do implicitly perform guided search.
4.1 Robot Modification
@@
There are many natural ways one might modify a robot description to obtain a closely related configuration space. A useful form of robot modification approximates the robot with simple geometries, such as bounding volumes, to speed up collision detection [29]. One of the most natural projections is to define $\mathcal{S}$ as workspace itself, ignoring the robot entirely and treating it as a fully controllable point robot in the environment [23, 15]. Some prior work ignores kinodynamic constraints on the robot [35], thereby yielding a guiding space where the robot is locally controllable. For chain-link robots, such as manipulators, a common guiding space involves focusing on a subset of the degrees of freedom [32]. Such methods often use heuristics in deciding which subsets to consider, e.g. focus on the first joints of a manipulator. Finally, some work attempts to learn a subspace of the robot’s degrees of freedom, e.g., using statistical projections such as PCA [33].
@@
All the above robot modification methods produce a search space as opposed to a guiding space. For example, workspace guidance methods may build a sparse graph in workspace, such as the medial axis [15]. Therefore there is often a secondary guided search that occurs in the modified robot space, e.g., the medial axis graph. This both highlights the hierarchical nature of guidance and helps isolate contributions - the medial axis is a useful structure for guiding search in workspace, and workspace is useful for guiding search in C-space.
4.2 Environment Modification
@@
Narrow passages are a known bottleneck for randomized sampling algorithms. As such, some work modifies obstacles in an attempt to widen passages [7, 39]. Lazy planning methods [11, 21] also perform a type of environment modification by ignoring constraints during initial planning, and incorporating them later as needed. When removing constraints, paths in $\mathcal{S}$ are not necessarily valid in $\mathcal{C}$, and thus often such methods iteratively or hierarchically compute a guiding space (e.g., the initial lazy graph), use $h$ to produce samples near the unvalidated path, and then repeat with a new guiding space informed by the recently acquired samples (e.g., remove edges found to be invalid).
In contrast to the removal of environment constraints, adding constraints is less common, but can be done to simplify geometries and therefore collision detection [19]. Moreover, paths found when adding constraints are always valid in the original search space. In other words, if we take $h$ to be the cost of paths in the modified space, constraint removal provides optimistic lower bounds whereas constraint addition produces conservative upper bounds.
4.3 Experience-Based Guidance
The most common version of experience-based guidance assumes the dataset contains a variety of environments all for the same robot. Consequently, database queries often consist of evaluating the degree to which paths in the dataset violate discovered constraints in the current motion planning instance [8]. While some methods only query using the task endpoints [14], more recent approaches propose to query (or filter) the database using local task features derived from workspace occupancy [13]. Another approach projects trajectories onto a lower dimensional abstract space via auto-encoding [25] or metrics like dynamic time warping [1], which improves database query. Some methods perform multiple queries, where the guiding space effectively becomes the composite space of multiple independent motion planning problems from the dataset, e.g., through gaussian mixture models [13].
@@ 4.3 Experience-Based Guidance
In all learning based approaches a major challenge is gathering appropriate data, leading to methods which simplify the problem by learning local guidance through learned local planners [17]. Some work [6] proposes to gather appropriate data for training by using SBMP.
5 Evaluation metric
Often we evaluate sampling based motion planning methods using algorithm level holistic metrics such as running time or number of samples. While minimizing such metrics is a reasonable end goal, they can be sensitive to implementation details, and focusing on them can obscure which aspect of a proposed contribution is responsible for the results. In the context of guided motion planning, we now present a new method to evaluate quality of guidance. We also note that prior work has considered alternative metrics for evaluating samples in motion planning, but in the context of giving a roadmap certain properties rather than guidance for a tree-based method [34, 12].
Tree-based search algorithms first select a node on the tree then expand the selected node to a new child. We focus on evaluating node selection as opposed to node expansion, and note that evaluating node expansion can be done by adding all possible (or a random sampling of) children to the tree, and using the evaluation of the best child according to our node selection evaluation.
5.1 Sampling Efficiency
Figure 2: A simple 2D example of the components of sampling efficiency: efficiency and (sub)-optimality. The nodes are colored on a green (best) to red (worst) scale. (a) and (b) show the efficiency score of each node, which corresponds to the work remaining. (c) and (d) show the suboptimality scores of the nodes, which captures the deviation of each node from the optimal path. Notice that the value of a node is relative to the other nodes, e.g., adding a new sample from (a) to (b) changes the value of all others. Notice that suboptimality depends on the tree structure, e.g., the rewiring between (c) and (d) changes the values of the nodes. Finally, despite depicting identical trees, (b) and (c) show how efficiency and suboptimality produce very different distributions over the nodes.
@@ 5.1 Sampling Efficiency
Intuitively, we aim to measure the difference between these two distributions. While the KL-Divergence provides such a measure, it can be dominated by small regions in space where the two distributions have the largest difference (e.g., when one distribution assigns zero probability to a region in which the other is non-zero). In sampling based motion planning we often aren’t worried about a few bad samples, so long as good ones are made with high enough probability so the bounded and symmetric Jensen-Shannon Divergence metric provides a smoother comparison, i.e.,
@@
In shortest path problems, we are interested in (efficiently) computing a boundedly optimal $(s,t)$ path in $\mathcal{C}$. Let $\delta_{v}$ be the suboptimality of $v\in T$, $\delta_{v}=d_{T}(s,v)+d_{\mathcal{C}}(v,t)-d_{\mathcal{C}}(s,t)>0$, where $d_{T}$ represents the distance between nodes on the tree and $d_{\mathcal{C}}$ is the true shortest path between configurations in $\mathcal{C}$. Now let $\tau_{v}$ be the work remaining for a search starting at $v$, $\tau_{v}=d_{\mathcal{C}}(v,t)$. Figure 2 shows an intuitive example of the relationship between these quantities. We propose to define $Q$ with respect to $\tau_{v}$ and $\delta_{v}$:
$Q_{T}(v)\propto\exp(-\frac{\delta_{v}}{\delta})\exp(-\frac{\tau_{v}}{\tau})=\frac{e^{-(\delta_{v}/\delta+\tau_{v}/\tau)}}{Z},$
$Z=\sum_{x}e^{-(\delta_{x}/\delta+\tau_{x}/\tau)}$
@@
We refer the reader to [4] for an in depth discussion of a metric to evaluate the quality of a sample for ST-connectivity problems in $C$-space; where measuring progress for the decision problem is not as straightforward as computing the cost-to-go.
5.2 Axes for Evaluating Guidance
The metric described above measures the quality of each individual node expansion step of the guided search algorithm, but there are other options.
5.2.1 Environment level
5.2.2 Sample level
In other words, single-step evaluation considers distributions over trees $T$, whereas multiple-sample evaluation considers distributions over $\mathcal{C}$, implicitly described by the distribution of valid trees produced by an algorithm. We refer the reader to [5] for an in depth discussion in this setting, and here summarize that in multiple-sample evaluation our metric reduces to measuring the entropy of the observed sampling distribution, where lower entropy means better guidance.
Figure 4: Example guiding spaces of each algorithm. We display the starts and goals of the rectangular robot for (a) RRT, (b) LazyPRM, and (d) PathDatabase to indicate that these guiding spaces are in C-space, whereas (c) MedialAxis is in workspace. (a) For RRT in SimplePassage, we show the Voronoi decomposition. (b) For LazyPRM in Cup, we show the unvalidated roadmap. (c) For MedialAxis in Trap, we show the workspace skeleton. (d) For PathDatabase in RandomSimplePassage, we display the paths from the database that pass near the goal.
@@ 5.2.2 Sample level
6 Refactoring Existing Algorithms
@@
Existing algorithms such as those cited in Section 4 are not implemented in the form of Algorithm 1, rather each uses their respective guiding space in a different way. In the next section we conduct experiments on four guided search algorithms, RRT [30], LazyPRM [11], DR-RRT [15], and Lightning [8]. Figure 5 shows examples of trees produced by these algorithms for solving the tasks depicted in Figure 3, and Figure 4 visualizes their guiding spaces.
@@
LazyPRM (environment modification) builds a probabilistic roadmap without performing collision detection with the environment (or only doing so for vertices but not edges). Our implementation grows a tree using the length of the shortest path on this lazy graph as guidance for selecting which node to expand on the tree. When an invalid sample is found in C-space, we pass this information back to the guiding space (the lazy graph) by deleting nearby vertices/edges.
@@
DR-RRT (robot modification) uses the medial axis skeleton computed in workspace to guide the robot. The original proposed implementation uses the edges of this skeleton to define “dynamic regions” along which to sample, and does so by creating balls around these regions and projecting back into C-space. For robots whose workspace configuration is a subset of the dimensions of their C-space this is easy, but for more complex C-space the general form of such a procedure involves an inverse-kinematics projection. Our new implementation, which we call MedialAxisGuidance is more straightforward - we project C-space nodes $v$ with forward-kinematics into workspace, then project these workspace poses onto the nearest medial axis edge $e_{v}$. The corresponding guidance value is then the distance along the skeleton edges from the projected edge to the projection of the goal, $d(e_{v},e_{t})$. Whenever a node $v$ fails to expand due to obstacle collision we increase the weight of $e_{v}$, thereby reducing the chance that the guidance encourages exploration along that edge.
@@
Lightning (experience-based guidance) uses a database of paths from related motion planning problems to guide exploration. Namely, given a motion planning problem, a path that was produced by the most similar task (start and goal) is selected from the database, and then copied into the current C-space. Where the path is invalid the algorithm fixes the solution by running a bi-directional RRT from the two ends of the breakage. We highlight that this sub-problem can be just as difficult as the original, and moreover that the full power of the database is never used. Our new implementation, which we call PathDatabaseGuidance uses the length of the queried path as the guidance value. As with LazyPRM, failed sample information is passed down to the guiding space by filtering out paths that pass close to invalid regions. If no node can locally plan to any path, then we default to RRT until a new node is found from which there is a visible path. We note that this implementation is relatively similar to [1], but we query the database using task similarity as in [8] rather than dynamic time warping, and more importantly we actively filter the database using information from the search tree.
Figure 6: RRT, LazyPRM, MedialAxis, and PathDatabase (Lightning) evaluations on TrapCup environment. All plots are averages over 128 random seeds. When a seed finishes the search we consider its sampling efficiency to be 00 for all future samples (i.e., sampling the goal). Runtime in PathDatabase includes only search time and not database creation, and guidance quality is averaged over the entire run instead of shown per iteration as in the other plots.
7 Experiments
@@
We design environments intended to demonstrate both strengths and weaknesses of each guiding space mentioned in Section 6, using a rectangular mobile base (3-dimensional C-space) in a 2-dimensional workspace. SimplePassage has two rooms connected by a narrow passage, Cup has a narrow “C”-shaped obstacle that creates a dead-end for search, and finally Trap contains two passages where the first is not valid in C-space. The evaluations are done with a smoothed target distribution that is at least $\epsilon=0.0001$ at each node on the tree. In addition, $\tau_{v},\delta_{v}$ values were normalized and we set $\tau=\delta=0.1$.
@@
Next we test PathDatabaseGuidance on a dataset of RandomSimplePassages, an environment with a randomly placed narrow passage between two rooms. We show how guidance improves as dataset size increases, while runtime gets worse (as querying the database takes longer), showing how such traditional metrics can obscure properties of algorithms because they are implementation dependent. This matches the intuition mentioned in Section 3.3 regarding the trade-off between guidance quality and the difficulty of computing it.
7.1 Multi-Guidance Search
Finally, we create a hybrid environment TrapCup which combines features from both the Trap and Cup environments: first the robot must avoid a passage that is valid in workspace but not C-space, then it must avoid a dead-end region. We now design a hybrid guiding space that is a combination of LazyPRM and MedialAxis. In other words, it creates both guiding spaces (i.e., the cross product) and uses a rule to determine how to select or combine the heuristics from the two. In the experiments shown in Figure 7 we simply swap the source of the guidance each time the search collides with an obstacle. Intuitively, each form of guidance faces certain local minima and by using multiple guiding spaces we can quickly escape those minima by passing control to another guiding space. Notice that the individual guidance plots for each guiding space indicate that these local minima are distinct (e.g., the curves have different peaks), and thus we see the hybrid guiding space out-performs its parts. Moreover, using number of samples alone as a metric would obscure this fact since MedialAxis and LazyPRM take about the same number of samples to find the goal. These results parallel what one would expect from running a multi-heuristic A* search [2] as opposed to single heuristic search, and highlights the value of framing sampling based motion planning as guided search; we now have a structured way of combining different algorithmic ideas from the literature as multiple guiding spaces, whereas the original implementations are not so conducive to modularization.
8 Conclusion